A square based boxed with an open top into be made from a square piece of cardboard that has side length 100 cm.The sides of the box are formed when four congruent square corner pieces are removed.The height. of the box to be formed is represented by the value x. Determine an expression for the area of cardboard used to make the box. ? If the surface area of the box is to be 6400 cm square,find the height of the box?
Answer by mananth(16946)

The side of the square you cut will be the height of the box.
you cut a square of side x
so area of aquare = x^2
you cut 4 squares
so area of 4 squares = 4x^2
The total area of the cardboard = 100 * 100 sq, cm
Area required for the box = total area - area of the cut squares
100^2-4x^2 =6400
4x^2=10000-6400
4x^2=3600
/4
x^2= 900
x= 30 taking the positive value